OpenStudy (anonymous):

HELP what is the phase shift of 3sin(x/2 + pi)

OpenStudy (anonymous):

pi

OpenStudy (anonymous):

I guess

OpenStudy (anonymous):

how did you get that??

OpenStudy (anonymous):

see pi inside sine function

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

that's called phase shift

OpenStudy (anonymous):

or phase change

OpenStudy (anonymous):

the phase shift...isnt its -c/b so wouldnt it be -pi/1??

ganeshie8 (ganeshie8):

Asin(Bx - C) phase shift = C/B

OpenStudy (anonymous):

wait...istn it -C/B

ganeshie8 (ganeshie8):

yes you're right ! for, Asin(Bx + C) phase shift = -C/B

ganeshie8 (ganeshie8):

3sin(x/2 + pi) here B = 1/2, C = pi just plugin :)

OpenStudy (anonymous):

okay so since its 3sin(x/2 + pi)

OpenStudy (anonymous):

1/2 / pi?

OpenStudy (anonymous):

oh other way around i meant!

ganeshie8 (ganeshie8):

-C/B = -pi/1/2 = -2pi

OpenStudy (anonymous):

yes!! thank you!!

ganeshie8 (ganeshie8):

np.. just remember that, since coefficient of x is a fraction, here, its period is more. and hence phase difference also will be more than the the constant term(pi here)

OpenStudy (anonymous):

.33 i think